Determine the IRR, to the nearets 1/10 percent, for an equipment that costs $250,000 and would provide positive cash flows of $100,000, $90,000, $80,000, and $20,000 at the end of each year for the next four years.

Complete the problem by hand using linear interpolation (show all calculations used, not the excel document) and also solve using Excel (show full document please). Hint: Start at 5%
Answer #1

IRR is rate at which Benefits equals cost

250000=100000/(1+irr)+90000/(1+irr)^2+80000/(1+irr)^3+20000/(1+irr)^4

Let 1/(1+irr)=x

20000x^4+80000^3+90000^2+100000x=250000

2x^4+8x^3+9x^2+10x=25

when irr=5% x=(1/1.05)=0.95

2(0.95)^4+8(0.95)^3+9(0.95)^2+10(0.95)=26.11

hence irr<5%

irr=6%

x=1/1.06

2(0.94)^4+8(0.94)^3+9(0.94)^2+10(0.94)=25.55

Bit closeer now irr=7.25%

x=(1/1.0725)

2(0.93)^4+8(0.93)^3+9(0.93)^2+10(0.93)=25.01

Hence irr is approximately 7.56%

By excel

Year
5%6%7.56%
110000095238.0952494339.6226492972.05607
29000081632.6530680099.679677794.22889
38000069107.0078867169.5426464290.57254
42000016454.049515841.8732614943.06679
Present Value Sum
262431.8557257450.7782249999.9999


answered by: AbschiedJuly